[Hallc_running] [New Logentry] Hall C status and near term plans (Dec. 18)

gaskelld at jlab.org gaskelld at jlab.org
Mon Dec 18 17:30:01 EST 2017

Logentry Text:
Beam was returned to Hall C last night with the aim of taking some large angle optics data in the HMS.  Ion chamber calibration issues again resulted in delays, but toward the end of owl/early day shift some data was accumulated on the "optics 1" target.

Today, we went to Controlled Access for work on the SHMS dipole. The dipole was deemed suitable for use for the remainder of the December, although with a current limit of 2100 A for now.

Tonight, we will take carbon elastic data with the SHMS at 7.5 degrees to check and verify the spectrometer tune. In parallel, data will be taken in the HMS to check trigger-level PID thresholds and scintillator discriminators thresholds. The hydrogen target will be condensed starting tomorrow morning.

The amount of progress made tonight with the SHMS optics will determine the program for tomorrow and beyond.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3506582
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171218/2008cd12/attachment.html>

More information about the Hallc_running mailing list